40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog, Page 1 of 58  Not logged in ELOG logo
ID Date Author Type Category Subject
  2911   Fri May 20 16:52:19 2022 ranaDailyProgressNoise Budgetmode matching, tfs, noise

I think in this plot, the actual phase margin is ~45 deg, not 225 deg. It would also be nice to indicate the gain margin on this plot.

(https://python-control.readthedocs.io/en/latest/generated/control.margin.html)

Quote:
 

Attachment 3: Open Loop transfer functions

 

  Draft   Fri May 20 14:46:14 2022 JCUpdateVacuumSearching for Leaks on Statlin

[Jordan, Shruti, JC]

Jordan and I headed over from 40 meter with the HLD. We did find that majority of the screws were loose on the middle flange and this may have played a part. After about an hour, we have yet to find a leak. The pressure will not go below 3 mTorr and we suspect that this leak should be large. We decided to turn on the Pfeiffer Vacuum Station and the turbo was able to reach 1500 Hz. Although, the pressure still did not change.

Our next step is to let the pump run over the weekend to see what the ultimate pressure is after 48+ hours. We will attempt again next week.


In attempt to run the pump over the weekend, the Hi Cube Vacuum turned off on its own and popped up an E006 Error. After searching for this error, we found this signifies the Run-up Time for the pump was not met. The remedies for this includes a leak in the system. We will follow up on next week for more trials.

  2909   Thu May 19 14:28:29 2022 shrutiDailyProgressNoise Budgetmode matching, tfs, noise

The present configuration: mode-matching less than 50%,

south laser with LIGO custom current driver [4,50 on knob] with ~ 4.5 mW of power entering cavity,

finesse ~ 9000 (=2pi/T_IC) with input coupler T=700 ppm

Attachment 1: While locked with gain setting 7.16 on knob, I measured the noise spectra.

Attachment 2: This is the PDH error signal which requires LO-RF phase re-adjustment that I realized only after measuring all the transfer functions and noises. I could not calibrate the noise with this.

Attachment 3: Open Loop transfer functions

We will have to repeat the above, after adjusting the PDH demod phase.

 

Other To Do:

  1. Replace mount on input coupler (the newport vertical mount is very close to clipping both the incident and reflected beams)
  2. Change mode-matching telescope to one in elog 2904.  We were not able to improve mode-matching beyond 50% with our previous telescope
  3. Add the plastic box with rubber base?
  4. Switch to locking with both teraxion (high power) and the south laser
Attachment 1: Spectra20220519.pdf
Spectra20220519.pdf
Attachment 2: PDHsignal.pdf
PDHsignal.pdf
Attachment 3: LoopTFs.pdf
LoopTFs.pdf
  2907   Thu May 19 10:32:35 2022 ranaDailyProgressPSOMAnew PSOMA cavity configuration

https://i.pinimg.com/originals/35/87/a6/3587a6ab4d6973b87d4756e5a3858791.gif

Quote:

I used the LB servo's input offset to null the current drive (with both inputs disconnected). Also moved the PDH error signal from -B to +A input on the LB servo. Finally, scanned around the temperature to maximize flashing on the reflection signal. WIth these adjustments, the cavity stably locks with LB box's gain knob set to ~7.65.

 

  2906   Thu May 19 10:16:04 2022 aaronDailyProgressNoise Budgetmode matching

I locked the cavity and added a 90-10 BS after the MC2 cantilever mirror to send transmission to both a DC PD and camera (90% reflection going to camera). While adjusting optics, the cavity remains locked (and even remains locked if the LB servo is turned off).

I improved the mode matching marginally by maximizing the ratio of transmitted to reflected light... but on introducing offsets to null the DC light level, the temperature control loop engaged with the cavity unlocked. Need to check its enable logic.

  2905   Wed May 18 19:32:57 2022 aaronDailyProgressPSOMAnew PSOMA cavity configuration

I used the LB servo's input offset to null the current drive (with both inputs disconnected). Also moved the PDH error signal from -B to +A input on the LB servo. Finally, scanned around the temperature to maximize flashing on the reflection signal. WIth these adjustments, the cavity stably locks with LB box's gain knob set to ~7.65.

  2904   Wed May 18 13:39:00 2022 shrutiDailyProgressPSOMAnew PSOMA cavity configuration

Here are the calculations for the mode mismatch between the horizontal (tangential) and vertical (sagittal) directions for when the curved mirror is being used as the input coupler.

Attachment 1 shows the mismatch (calculated at the point with the smallest beam) with the angle of incidence on the curved optic and Attachment 2 shows the ideal minimal beam waist for the two directions. The effective radius of curvature seen for a beam at \theta angle of incidence is R \cos(\theta) in the tangential direction and R/ \cos(\theta) in the sagittal direction.

Attachments 3 and 4 are two new mode-matching solutions considering the new location of the curved mirror, but still only using spherical (not cylindrical) lenses without correcting for the astigmatism.

Attachment 1: Mismatch.pdf
Mismatch.pdf
Attachment 2: BeamRadius.pdf
BeamRadius.pdf
Attachment 3: MMsol1.pdf
MMsol1.pdf
Attachment 4: MMsol2.pdf
MMsol2.pdf
  2903   Wed May 18 12:56:35 2022 aaronDailyProgressPSOMAnew PSOMA cavity configuration

I added the South EOM into the fiber path for the Teraxion laser. With the modulation on, I now see the PDH error signal while the transmission and reflection are flashing. I connected the LB servo box output to the mod in port of the Teraxion laser using a long BNC and BNC-to-SMA adapter.

After adjusting the input alignment to maximize flashes on the transmission PD, I moved the transmission camera from MC3 to MC2; the beam fills more than the full field of view of the camera after MC3, because MC3 is the curved mirror. I was unable to locate the transmitted beam through MC3. With the camera at MC2, I can easily located the flashing transmission beam.

I adjusted the input steering and LB servo gain until I lock on to a 00 mode for 10s of seconds.

  2902   Wed May 18 11:19:34 2022 shrutiDailyProgressPSOMAnew PSOMA cavity configuration

With the current high power level (~40 mW) entering the cavity, I was able to align so that I could visibly observe one round trip along the cavity.

The vertical mount for the input coupler/ MC1/ curved mirror has the beam entering and exiting very close to its edges for the 45 deg angle of incidence.

I also began to see dips in the reflection and peaks in the transmission from the cantilever (the same forest of peaks as before). So, very close. Since this is the teraxion laser, I was not sure how to modulate it while aligning to get max mode-matching.

  2901   Tue May 17 14:28:09 2022 aaronDailyProgressPSOMAnew PSOMA cavity configuration

Shruti tells us that our 1m curved mirror operated at 45 degree incidence as MC1 will introduce a 6-7% power mismatch due to astigmatism. We can accept that for now, so I'm making the swap discussed above.

Here is the new configuration and measurements of the transmission through each optic

Position Optic Power incident Power transmitted Transmissivity (T)
MC1 1m FL curved mirror      
MC2 cantilever 40 mW 20 uW 500 ppm
MC3 coastline high reflector 40 mW 1.6 < 50 ppm

I mounted MC1 and MC3 in the top-adjusting mounts from Newport. They seem pretty good (take up very little space on the breadboard despite having pitch and yaw adjustment). However, despite aligning the beam with MC1 removed several times, I wasn't able to complete the alignment with MC1 in place due to clipping on the mount and the beam wandering off of the cantilever (MC2, which is a much smaller aperture mirror than previously).

  2900   Fri May 13 15:02:29 2022 ranaDailyProgressPSOMAnew PSOMA cavity configuration

An added comment to Aaron's entry below:

Let us use the naming convention which is commonly used in LIGO:

MC1 = input coupler

MC2 = back mirror at vertex

MC3 = the other 45 deg mirror

In the previous config, the curved mirror was MC2, and the canitlever was MC3, and the Coastline high reflector with T = 25 ppm was MC1.

New config:

MC1 = curved Si mirror

MC2 = cantilever

MC3 = Coastline HR (T = 25 ppm)

The new issues is that the 45 deg incidence angle will create some astigmatism in the cavity. We (Shruti) need to compute how much of a mode-matching hit this will be. I expect it will be small, so I suggest to make the change to the cavity and start locking while Shruti computes the mode mismatch. Since the mismatch is proportional to ((delta w)/w)^2, we should adjust the mode matching to get a circular waist halfway between the vertical and horionztal waists of our astigmatic cavity. IF the mismatch is going to be more than 15%, we should buy some cylindrical lenses.

  2899   Fri May 13 12:57:30 2022 aaronLab InfrastructureVacuumleak checking cryo cantilevers vacuum

Jordan will soon bring over the leak checker to debug our Statlin cryostat.

I powered up the vacuum gauge and connected it to the acromag. I confirmed that the slow channel X1:OMA-ERC_PRESSURE_UTORR is still recording the vacuum pressure.

I also connected the vacuum pump to our cryostat and powered on the pumping station (but did not pump down). We eventually need a solution for routing the vacuum tube through the enclosure, or mounting the turbo inside the enclosure.

I also made a new epics channel on cominaux that calculates the log of the vacuum pressure (in torr). However, despite following Chris' steps I wasn't able to get the channel logging to frames.

  2898   Thu May 12 11:22:44 2022 aaronDailyProgressPSOMAaligning

[shruti, aaron]

We switched to the teraxion laser today for a higher power to be able to align more easily.

Initially, we hooked up the output of the laser head to the Faraday isolator on the Cryo cavs table and then replaced the input to the "South EOM" with this light. With the EOM we saw 15 mW after the free-space coupler. After removing the EOM from the path we saw almost 40 mW. Transmitted through the input coupler, we would expect 2.4 microW but the max we could measure on the power meter was .8 microW

[+rana]

We measured the transmissivity of the cavity mirrors, and found the cavity was far too undercoupled (measured transmissivities attached). The cantilever is far too transmissive at 45 degrees (0.7%) for a high finesse cavity. By measuring the mirrors at both normal and 45 deg incidence, we identified the attached alternative configuration for the cavity to try next time.

Attachment 1: WorkNB_Y3Q3_2.pdf
WorkNB_Y3Q3_2.pdf WorkNB_Y3Q3_2.pdf WorkNB_Y3Q3_2.pdf
  2897   Wed May 11 14:11:15 2022 aaronDailyProgressPSOMAaligning

Thanks, that description was very helpful for understanding the state of the cavity as soon as I walked in.

I measured the laser power at various points along the beam path, and found there's still quite a bit of scatter -- most of the power transmitted through the cavity is not blocked if I block the prompt transmission through the cavity. I also had trouble picking up the single pass beam on the TRANS PD.

It would be easier if we had more incident power, and a wider clear aperture for the input coupler. Therefore, I bypassed the 50-50 fiber BS that was mixing N and S lasers, and instead added a patch cable between the S EOM and the beam launch. The S laser now provides 4 mW (rather than 1.8 mW) out of the fiber launch. It could provide up to 7 mW, but the REFL PD can only handle 5 mW; we aren't well enough aligned for it to matter yet, but I don't want to risk frying the PD.

I also found a Thorlabs H45 adapter in our optics closet. We ordered a few of these from Thorlabs because they can be added to a kinematic mount to avoid the beam clipping on adjustment knobs. I replaced the optic that was in the H45 (unlabelled) with our high reflector (45 degree 99.994% reflectivity, from coastline)

I then aligned with the steps above. Modifications

  • I pulled out a clean beam dump and place it after the input coupling mirror to avoid scatter from downstream optics

I aligned the beam path to the REFL PD without the input coupler, but haven't completed one round trip with the input coupler in place.

  2896   Wed May 11 12:42:04 2022 shrutiDailyProgressPSOMAaligning

Initially I began with the procedure in elog 2895, but realized that the height of the beam was not parallel to the surface of the table so I removed the input coupler and made sure the height of the beam was at 4 inches everywhere.

I also got the beam to be incident at the center of the curved mirror and then to the center of the cantilever miror. After that, I checked that the height of the beam after the cantilever was also roughly at 4 inches up to the steering mirror to the PD 1811 (this was still without the input coupler).

When I added the input coupler, I adjusted its vertical direction and then got the beam to the PD.

Using the diagram in Attachment 1 as a guide, assuming the beam is perfectly horizontal (parallel to the table surface), I played around with the horizontal alignment.

Attachment 1: IMG_D3C4C21EF227-1.jpeg
IMG_D3C4C21EF227-1.jpeg
  2895   Tue May 10 14:56:42 2022 aaronDailyProgressOpticsmeasured scatter for a while, but here's the cavity alignment procedure

Continuing with alignment.

There are several spurious oscillations (10s of Hz, few kHz, few MHz) on the REFL PD's DC signal displayed on the oscilloscope. Disconnecting the BNC from the REFL DC MON to our ADC removes the oscillation, so I've done that to help with alignment.

I also noticed I was seeing ~50 nW of light on the power meter due to the table lamp I was using in lieu of room lights. After several minutes of aligning, I noticed the maximum achievable 'beam' power decreased monotonically along one round trip in the cavity (as if it were widely diverging) and that the shadow of my arm could affect the power meter readout. I turned off the lamp and will align in the dark.

Here's what I'm doing to align:

  1. Use a viewcard and power meter to ensure the beam enters and exits the cavity without clipping, checking just before and after the input coupling mirror.
  2. Adjust the output steering mirror to center the beam on the REFL PD
  3. Place the power meter just before the curved mirror (second cavity mirror encountered by the beam), and adjust the input alignment to maximize power on the power meter. Adjust the output steering mirror to ensure the beam is still incident on the REFL PD, and check for clipping on the output.
  4. Place the power meter just before the cantilever mirror (third and final cavity mirror encountered by the beam), and adjust the curved mirror to maximize power reported. I'm seeing 50-70 nW.
  5. Place the power meter just after the cantilever mirror, and adjust the curved mirror to ensure the beam reflects off the cantilever mirror.
  6. Place the power meter behind the cantilever mirror and check monitor the transmitted power. If the beam makes multiple round trips, the transmitted power will be more than that measured inside the cavity. Adjust the curved mirror and input coupling mirror to maximize power on the cavity, then adjust the output steering mirror to center the reflected beam on REFL PD again.
    • placing the power meter farther back from the cantilever mirror forces a tighter alignment, because diverging beams will not all reach the power meter.
    • when placing the power meter, block the cavity from making a full round trip, or you'll only pick up the beam that's missing the cantielver
    • The power should increase in steps according to the transmitted light from one round trip (so P_in * T_input * T_cantilever). This is especially true when observing the second round trip. If it jumps with a step size comparable to the off-resonance circulating power (P_in * T_input), the beam is missing the cantilever. Move the power meter farther away from the mirror.
  7. Eventually, switch over to monitoring transmission on a DC photodiode rather than the power meter. It's faster and more sensitive. I adjusted the height and orientation of the TRANS PD until the signal was maximized, then continued adjusting input and curved cavity mirrors.
    • This all seemed fine, until I went to adjust the output steering to make sure the beam still hits REFL PD. For some reason moving the output steering mirror also changes the oscilloscope trace for the TRANS MON signal. Just to be clear about how weird this is:

Diversion

  • Blocking the TRANS beam sends TRANS MON trace to 0 as expected... but blocking the REFL beam ALSO causes the TRANS MON trace to go to 0!
  • Blocking both beams is the same as blocking one beam.
  • This occurs even if I disconnect the REFL PD from power... it makes me suspicious of ghost beams, but occurs even when I place a viewcard directly in front of the (unpowered) REFL PD...
  • In fact, I observe this behavior even when the REFL PD is disconnected from power, and its DC and AC outputs are disconnected! I'm 100% sure I'm not clipping the beam elsewhere. It happens when I block the beam with a viewcard or another, less shiny material (finger, paper).
  • ah, Shruti points out that I'm seeing the backreflection. I thought I was safe because the PD is at an angle relative to the penultimate mirror, but the lens curves the beam and the TRANS MON PD is large. Well attached is some ghostly behavior.
  • I rotated the REFL PD to separate the backreflected beam, and adjusted the lens to maximize signal on the PD. I was able to increase the reflected power substantially -- the REFL PD's DC MON reports 1.86 V, compared to 1.75 mW measured on the power meter. The PD responsivity is 1 A/W and its DC gain is 10 V/mA, so we're still losing a lot of light.

Unfortunately, it appears I was aligning the cavity to the backreflected beam this whole time. So, I'm repeating the above steps.

  • Oddly, this time I'm seeing 500+ nW on the power meter placed inside the cavity. I'm expecting more like 100 nW, since I have 1.8 mW incident and Johannes measured 99.994% HR reflectivity. We could add a waveplate to rotate the polarization, but I'll proceed for now.

Ah, quickly learned that the scatter off of the edge of optics mounts can be O(100 nW - 1 uW). Enough of this for today. Enjoy the video, book my magic show for your birthday party.

Attachment 1: trim.5617A4E4-1A7A-49C2-B66B-9D65F3013712.MOV
  2894   Tue May 10 14:05:56 2022 shrutiDailyProgressOpticsReplaced the cavity input coupler with 99.994% R optic for higher finesse cavity

I switched the optic so the HR surface is inside the cavity. Still working on the alignment.

  2893   Thu May 5 17:09:06 2022 aaron DailyProgressPSOMAloop TFs and noise spectra

[shruti, aaron]

For sanity purposes, we also plotted the frequency noise spectrum as recorded by the control point. In this case, we undid both the cavity pole and the LB box transfer function to arrive at the frequency noise at the cavity input, then undid the loop suppression in the same manner as for the error mon traces. The plot showing both the control point estimate and error point estimate of the out of loop frequency noise is attached.

There is some discrepancy at high frequency, but the 100 Hz to 10 kHz band has good agreement.

Our intention originally was to inject a tone at ~kHz into the loop, then independently measure the resulting frequency modulation using the PDH error signal and another frequency standard... however, we have independent frequency standard for comparison. We have not locked both lasers to the cavity, so the beat note wanders too much to measure its frequency with the delay line box or phasemeter.

Perhaps another way make this check would be to refer to the cavity input in both directions around the loop: by undoing the cavity pole and PDH gain, vs by propagating forward through the current driver and laser responsivities. I did so in attachment 2, but only applied a uniform gain for the diode and current driver responses; I believe Shruti may have measured their transfer functions, so we could update this check to reflect the frequency dependent gain.

Attachment 1: LB_CTRL_comparison.pdf
LB_CTRL_comparison.pdf
Attachment 2: LB_CTRL_comparison.pdf
LB_CTRL_comparison.pdf
  2892   Thu May 5 17:01:40 2022 shrutiDailyProgressOpticsReplaced the cavity input coupler with 99.994% R optic for higher finesse cavity

[shruti, aaron]

We replaced the 90-10 BS used as the input coupler to the cavity with the Coastline optics HR mirror.

We did this by only removing the optic from the mount while keeping everything else in place (mount, post, knobs).

While doing this we realized that the HR surface had to be on the outside of the cavity in order to be able to see a reflected beam. The three knobs of the mount were blocking the beam otherwise.

Photos and videos of the swap are available on our google drive.

  2891   Thu May 5 14:09:42 2022 aaronDailyProgressPSOMAloop TFs and noise spectra

I noticed that the oscilloscope was still plugged in to a network switch, but the switch itself was not plugged in to the router. Once it was plugged in, I identified the scope's ip with nmap as 10.0.5.247 and grabbed the data with

python tek_tools.py 10.0.5.247 'ch3' ~/cryo_lab/data/PDH/cal 220502.csv
  2890   Wed May 4 15:52:56 2022 shruti DailyProgressPSOMAloop TFs and noise spectra

After some troubleshooting I realized that I was mistakenly measuring the PDH signal coming from the not-in-use north laser PDH mixer+LPF electronics box, so I changed that to the right one.

Now the LB1005 error monitor and PDH error point show pretty much the same noise spectrum [Attachment 1]. Input 1 (pink)- PDH, Input 2 (blue)- LB mon, Input 3 (green)- Ctrl. 

I re-measured the loop transfer functions and they are so much less noisy. Also seems to have a UGF of 170 kHz! [Attachments 2, 4]. Attachment 4 is the same as Attachment 2 and includes the UGF calculated from raw data and not just the simple fit. Something weird happened in the closed loop TF measured at the control point which did not effect the OLTF estimation. The y-scale is in dB for the magnitude.

The estimate of the unsuppressed noise is now much higher after accounting for the new transfer functions for the data from 03-May-22 measured at the LB mon point [Attachment 3]. Compare with earlier estimate.

Latest noise budget script with results.

 

Attachment 1: 1DFC2C2A-ABB4-41ED-9F2C-7A667BAA4B96.png
1DFC2C2A-ABB4-41ED-9F2C-7A667BAA4B96.png
Attachment 2: LoopTFs.pdf
LoopTFs.pdf
Attachment 3: CalibratedNoiseBudget.pdf
CalibratedNoiseBudget.pdf
Attachment 4: LoopTFs.pdf
LoopTFs.pdf
  2889   Tue May 3 16:34:19 2022 ChrisDailyProgressPSOMAcurrent driver connections and cantilever acoustic damping

I opened the LIGO current driver chassis for the north laser, to check its PCB revision. It is v5 and will need modifications similar to those applied to the south laser's driver.

Cymac ADC channels 8-11 (zero indexed) were connected to the DAQ monitor signals on the south laser's driver. New filter modules X1:OMA-SLD_CATHODE, X1:OMA-SLD_HF_IMON, X1:OMA-SLD_LF_IMON were added to the x1oma model for access to these signals. Only the Imon signals are useful at the moment; the cathode voltage monitor overflows the ADC range.

While updating the model, I added a CTRL subsystem, intended for slow offloading of the control signal onto the current driver's LF input and/or the laser temperature. (See attachments 1 and 2.) These offloading paths can be triggered by the TRANS_MON signal.

Cymac DAC channel 4 (zero indexed) was connected to the LF current input, and the loop was closed to suppress the control signal offset (settings in attachment 3).

The cavity transmission was found to be intermittently glitchy (see attachment 4). Glitches seem to get excited along with the cantilever motion, but sometimes vanish for long stretches of time. We should probably check that the laser is not near a mode hopping region, cavity HOMs are not near degeneracy, etc.

Finally, as an experiment in acoustically damping the cantilever, I connected DAC channel 5 (CTRL_TEMP output) to the stereo speakers through an SR560 (gain=1; stereo volume=10). With a bandpass around the cantilever resonance and some tweaking of gain and phase settings (attachment 5), it was possible to knock the 44.5 Hz peak down by a modest factor (attachment 6, damping=red, no damping=green reference trace). The second harmonic was well suppressed also, but there was little visible effect elsewhere in the spectrum.

MEDM screens still have to be updated for the model changes.

Attachment 1: x1oma.png
x1oma.png
Attachment 2: x1oma_ctrl.png
x1oma_ctrl.png
Attachment 3: ctrl_current.png
ctrl_current.png
Attachment 4: signals.png
signals.png
Attachment 5: ctrl_temp.png
ctrl_temp.png
Attachment 6: damp_dtt.png
damp_dtt.png
  2888   Tue May 3 15:34:15 2022 aaronDailyProgressPSOMAloop TFs and noise spectra

I updated the calibrated noise budget based on Shruti's data above.

Shruti: Attachment 2 is the same using the LB error mon point

Attachment 1: CalibratedNoiseBudget.pdf
CalibratedNoiseBudget.pdf
Attachment 2: CalibratedNoiseBudget.pdf
CalibratedNoiseBudget.pdf
  2887   Tue May 3 09:51:42 2022 shrutiDailyProgressPSOMAloop TFs and noise spectra

Loop transfer functions

[Attachment 1]

  • For excitation at the 'A' port of the LB box, I measured the LB monitor (right after excitation), PDH signal (after the excitation travels the loop), and the control signal (after the LB).
  • The phase coherence of the PDH signal was quite low for excitations below 130 mV. So I measured the spectra multiple times with different excitations and plotted all of them together - some traces are less noisy but in the new somewhat automated data acquistion scheme I plotted everything together.  Below 1 kHz I had to use over 200 mVpp. The higher excitation is probably because the new HF drive has a mA/V coeff 100x smaller (0.2 mA/V iirc) than the ITC502 (20 mA/V).
  • For the next measurement I would probably try to excite at a different location because increasing the excitation further here causes the loop to lose lock

PDH calibration

I tried getting the data from the oscilloscope using the same script as earlier but saw 'connection refused' or 'connection timed out' each time. I did verify with `nmap` that the address I was using was correct. Not sure what was happening. Therefore, for the noise calibration I used the previous values.

python tek_tools.py 10.0.0.249 3 ~/cryo_lab/data/PDH/cal 220502.csv

Noise spectra

Attachment 2 - calibrated using the spectra measured at the PDH signal location

Attachment 3 - the raw voltage noise traces at the PDH, LB monitor, and Control points

Attachment 4 - Comparison of the noise spectra. Dashed - using PDH point, Bold - Using LB error mon, Dotted - previously measured noise.

Attachment 5 - The forest of peaks around and above a kHz is also now not as visible in the spectra measured with diaggui with the cavity locked. We certainly did see it earlier, not sure if we have any data saved

Overall, the noise looks much lower than the previous measurement and does not seem to have the forest of peaks. We will further update this plot with the other terms in our budget shortly and try to make sense of it.

I will add the pomona box to get additional gain at the cantilever resonance to try to further dampen the noise at 45 Hz and the higher harmonics of the cantilever

 

Questions

  1. Why is there a discrepancy between what is measured at the LB error mon and the PDH error signal (at the output of the mixer+LPF)? There is only a unity gain buffer in between. Does this mean the input is saturated? [to test] (05-Mar-22: realized PDH error signal was not measured from the right port)
  2. What is the strange peak at 900 kHz in the spectra? Also seen in laser frequency noise [ref elog 2886]
Attachment 1: LoopModel.pdf
LoopModel.pdf
Attachment 2: Noise.pdf
Noise.pdf
Attachment 3: Spectra20220502.pdf
Spectra20220502.pdf
Attachment 4: Noise.pdf
Noise.pdf
Attachment 5: noisespectra1.pdf
noisespectra1.pdf
  2886   Fri Apr 29 11:21:51 2022 shrutiUpdatePSOMASwitch current driver to custom low noise

Using the delay line setup with a digital PID slow control on the Moku:Pro in multi-instrument mode, I first calibrated the delay line by offset locking and measuring the beat frequency with small offsets, and then measured the noise spectra at the delay line with no offset. The script I used to analysze the results is here.

Then I ran the noise calibration script and plotted this instead of the previously measured data [Attachment 1].

Attachment 2 contains a comparison of the old and new laser noises measured with the North-South beat. The north laser is still powered by the ITC510 which probably is why the forest of peaks still exist.

 

Attachment 1: CalibratedNoiseBudget.pdf
CalibratedNoiseBudget.pdf
Attachment 2: laser_noise_comparison.pdf
laser_noise_comparison.pdf
  2885   Thu Apr 28 14:44:28 2022 aaronUpdatePSOMASwitch current driver to custom low noise

[shruti, aaron]

PDH calibration

Later in the afternoon, we again measured the PDH error signal sweep driving with 2V from the function generator at 1.3 kHz. I saved the trace in cryo_lab/data/PDH/cal/220428_PDHcal.csv.

The best fit parameters are

  • Slope of the error signal is 0.33 V/MHz, or 2.99 MHz/V
  • Fitted cavity length is 30.1 cm, and the FSR is 998 MHz
  • Fitted transmission of the cavity is 1.0%
  • The relative RF and LO phase at the mixer is 198 degrees (nearly out of phase, off by 18 degrees)

Dark noise measurement

We blocked the beam and measured the noise on the PDH error signal using an SR785. We sent the PDH error signal from the error mon point on the LB servo box (same location we used for the calibration above) directly to the spectrum analyzer.

The new curve for photodiode dark noise is at 1 Hz/rtHz above a few kHz, and the curve lies below our previous estimate for the frequency noise of the laser -- progress!

Calibration pipeline

I'm making a file calpar.yml in our scripts/calibrate_noise directory. This file will hold the key variables for generating the noise budget, including file names and paths, cavity parameters, and which PDH calibration to use for various measurements.

Beat note

Shruti measured the frequency noise on the beat note in the same manner as above. More details and results to follow.

Shot noise

We didn't measure the shot noise, but we estimated to see the shot noise above the observed photodiode dark noise would require about 3 mW of power at 1550nm on the photodiode. We can then scale the shot noise by the actual power level with the cavity locked.

Attachment 1: fitted_full.pdf
fitted_full.pdf
Attachment 2: CalibratedNoiseBudget.pdf
CalibratedNoiseBudget.pdf
  2884   Thu Apr 28 11:19:24 2022 shrutiUpdatePSOMASwitch current driver to custom low noise

[chris, shruti, aaron]

 We turned down the current on the Thorlabs ITC502, rack mounted the current driver updated by Chris to D1200719 - v7, and used that to driver the south laser.

We adjusted the coarse knob to 4,90 to see roughly the same refl DC voltage as measured on the scope (which should translate to what I thought was ~135 mA from looking at the 'ANALOG OUT' on the ITC502 but seems to be <100 mA by Chris' measurement). The ITC 502 is still used as the temperature controller, the settings of which were not changed at this stage.

Then, we tried to lock the cavity. The ITC502 HF driver coefficient was 20 mA/V and the LIGO custom one is 0.2 mA/V. This required us to make few changes to the LB1005 settings and its output electronics. With a knob setting on the LB of 6,60 and nothing else in the output chain to the HF drive input of the driver (we removed the attenuator, 50 Ohm terminator that was teed, no pomona filter was used), we were able to lock the cavity for a very long time ( >5 min). The PDH signal noise when locked was ~150 mV pk-pk, dominated by the 45 Hz oscillations, relative to a pk-pk signal of 600 mV (Not quite sure of this value since the PDH signal was locked with a large offset  of ~-1 V in the output current and may have been at a different lock point than what we saw -- would need the slow temperature control to zero this).

 

 

 

  2883   Thu Apr 28 06:12:14 2022 ChrisElectronicsPSOMACurrent driver repair/upgrade

One of the D1200719 current driver boards (S1500270) was in the EE shop for repair after some unexplained current fluctuations were seen.

Previously I measured its output at different knob settings, and saw saturation when driving a 25 Ω load. The saturation happens when the low frequency current source block reaches its voltage limit. (This wasn’t seen in Zach’s previous test, probably because a lower-impedance load was used.)

The board has now been upgraded from v4 to v7, to fix high frequency oscillations and crossover issues. An annotated schematic is uploaded to the board’s DCC traveler.

A few tests were made to check the basic functionality:

  • A 25 Ω dummy load was connected for testing (note: this component must be rated for high power dissipation)
  • The unit’s nominal current draw from the supplies was about 200 mA
  • The response of the coarse and fine adjust knobs was measured. (Saturation occurs above a coarse knob setting of 7, when driving the 25 Ω load)
  • Noise spectra were measured on the SR780 analyzer at the HF and total current monitor outputs, with a coarse knob setting of 4. (Both were nearly white, ~10 nV/rtHz HF and ~20 nV/rtHz total)
  • A wideband transfer function from HF modulation input to HF monitor output was measured, using 4395A analyzer + ZSC-2-1 splitter + HAT-3+ attenuator on the differential input (which was driven single-ended)
  • The transfer function from the DAC input to the total current monitor output was checked (showing the expected 40 dB/decade filtering above 10 Hz)
Attachment 1: knobs.pdf
knobs.pdf
Attachment 2: spectra.pdf
spectra.pdf
Attachment 3: AG4395A_25-04-2022_170840.pdf
AG4395A_25-04-2022_170840.pdf
Attachment 4: data.zip
  2882   Tue Apr 26 16:18:11 2022 aaronNoise HuntingNoise Budgetcalibrating PDH error signal

I'm grabbing the pdh error signal from the oscilloscope over ethernet (labutils/tektronix/tek-grabber.py), and will calibrate the PDH error signal.

Just from examining the oscilloscope trace, the sideband zero crossings appear to be about 1.1 ms apart, while the DC resonance error signal is 708 mV tall and 56 us wide (measured with the cursor at the peaks). I think the slope of the error signal passing through resonance is related to the peak-to-peak measurements by

\frac{\partial \epsilon}{\partial f} = 2 \frac{\Delta V_\mathrm{pkpk}}{\Delta t_\mathrm{pkpk}} \frac{\Delta t_\mathrm{sideband}}{\Delta f_{\mathrm{sideband}}}

This is based on Eric Black's introduction to PDH stabilization, namely

D\equiv -\frac{8\sqrt{P_cP_s}}{\delta \nu} in equation 4.2, where D is the frequency discriminant and \delta \nu the linewidth.

And from Black's Fig 7, the PDH error signal's peak-to-peak height is 4\sqrt{P_cP_s}

Since the sideband frequency is 33.59 MHz, this implies the slope of the error signal is 0.41 V / MHz, or 2.4 MHz / V. I might be off by a factor of \sqrt{2}, depending on the definition of the linewidth in Black. Anyway I'll check this against the fit from PDH_calibrate.ipynb. In the end, looks like the by-eye estimate nearly agrees with the fitted result.

In PDH_calibrate.ipynb, I added a parameter to characterize the relative phase between the LO and RF signals. I think I've implemented this correctly, and find a 13 degree detuning of the LO from the RF signal at the mixer.

[stymied again by the lab ethernet and other distractions. Not seeing the oscilloscope show up on nmap, possibly just a bad cable. anyway will pick it up tomorrow morning]

update 4/27/22:

Simple debugging on the oscilloscope frontend let me grab the PDH trace via ethernet.

I tinkered around with PDH_calibrate to track down various factors of few and pi. Since the script fits the entire PDH error signal (swept through all three resonances), I can extract a couple of extra parameters that should be measured indepently for sanity:

  • The round trip cavity transmission (really 1-R) is 0.01, implying a finesse of 308.
  • The 'best fit' cavity length is 32 cm and the FSR 944 MHz

The best fit slope of the error signal is -0.46 V/MHz or 2.2 MHz/V

Attachment 1: fitted_full.pdf
fitted_full.pdf
  2879   Tue Apr 26 14:28:51 2022 aaronThings to BuyOpticshigh reflectors in lab

We have been wondering about the availability of high reflecting optics for the PSOMA cavity, especially the input / output coupler at 45 degrees incidence. I located about 25 of the coastline mirrors measured by Johannes in 2018 with an HR reflectivity of about 99.994% and AR reflectivity 0.04%. Along with the 1m, 180 ppm curved end mirror we have been using, the coastline mirrors should allow us to achieve a cavity finesse in excess of 10,000. I suggest we swap in these mirrors as soon as the noise budget is at all sensible.

If we require high reflectors at normal incidence, we may need to place another order, or check the properties of some of the other custom optics on our shelf.

  2878   Mon Apr 25 13:54:03 2022 shrutiNoise HuntingNoise Budgetcalibrating PDH error signal

After checking that the cavity locks, around the same temperature and current set-point, sweeping the current resulted in something that looked closer to a PDH signa [Attachment 1]. I think adjusting the RF phase wrt the LO before the mixer, by changing the relative length of the cables (10 m ~ 1 wavelength at 30 MHz; the modulation is at 33.59 MHz) should get the signal to look more like a PDH demodulated signal with a larger linear range around resonance.

I adjusted the cables between the OCXO and EOM and did manage to recover a proper looking PDH signal [Attachment 3]. Given that there were not many available cables to play with it is probably still not optimal.

I also increased the gain on the LB1005 slightly from 5.7 to 5.9 knob setting value.

 

Attachment 1: IMG_1620.pdf
IMG_1620.pdf
Attachment 2: IMG_1619.pdf
IMG_1619.pdf
Attachment 3: IMG_1622.pdf
IMG_1622.pdf
Attachment 4: IMG_1623.pdf
IMG_1623.pdf
  2877   Thu Apr 21 18:13:36 2022 Ian MacMillanMiscEquipment LoanAccelerometer/seismometer loan Wilcoxon 731A

Returned. See this elog

Quote:

I borrowed a Wilcoxon Seismometer Model number 731A from the cryo lab. I am also going to borrow an amp/power source for it. Currently at the 40m

 

  2876   Wed Apr 20 15:48:07 2022 aaronNoise HuntingNoise Budgetcalibrating PDH error signal

I'm calibrating PDH error signal first to avoid recording data of 'dark noise' with no PDH signal like yesterday

I first locked the S laser to the cavity using the LB servo box in the usual configuration. I drove the 'sweep' input of the LB box with a 3.14 kHz triangle wave from a function generator, and monitor the PDH error signal at the 'error monitor' port of the LB box using an oscilloscope. I set the amplitude from the function generator at 1 Vpp, and tuned the sweep amplitude on the LB box until the sweep took the laser through both sideband resonance but not the +- FSR resonances. Then, I adjusted the timescale on the oscilloscope to capture one full sweep through both resonances, and triggered 'single shot' until I captured all three resonances.

I didn't get a clean error signal -- instead I'm always seeing a flat region as the DC frequency sweeps through resonance. Attachment 1 shows the PDH error signal in purple (ch3), refl in blue (ch 2), and trans in yellow (ch1).

I thought maybe the LB servo was holding the frequency on resonance (despite the servo being disengaged), but turning the servo gain down to 0 does not affect the error signal trace. Maybe I should watch the error signal before the unity gain buffer amplifier that precedes the LB box's error monitor? Probably someone (maybe myself in a previous elog) has seen this, so I'll ask around.

Attachment 1: 9B21DA6F-14BA-453D-93BB-C6C8803CBCA0.jpeg
9B21DA6F-14BA-453D-93BB-C6C8803CBCA0.jpeg
  2875   Wed Apr 20 14:27:41 2022 aaronLab InfrastructureGeneralstarting up after power outage

I'm turning on the lab electronics after the planned outage this morning.

  • Turned on following electronics.
    • cymac1, gaston, spirou, and cominaux
    • Sorensen DC supplies, and the GPS-30300 supplying 8 VDC. With all other electronics listed below powered on, the Sorensens supply +18 V with 2.7 A, and -18 V with 1.0 A.
    • AA, AI, and acromag chassis.
    • PSOMA rack Rio laser drivers, and engaged both TEC
    • PSOMA REFL, TRANS, and beat note photodiodes
    • PSOMA table oscilloscopes and TV
    • moku
    • PSOMA enclosure HEPA FFU (the one unit that fits on the enclosure)
  • Ran standard cymac startup script
    • rtreset
  • Confirmed status of digital system
    • systemctl status on cominaux is 'running'
    • testpoints on the realtime system are green, except the IOP timing indicator (red as usual) and the DAQ_DETAIL GPS indicators (white boxed as usual). See attachment 1.
    • The temperature and particle count channels were dead... in fact they haven't been recording for almost 6 months no.
      • I restarted rts-edc on cymac1, but the channels still read 0 despite the digital display on the particle counter reporting nonzero values
      • I noticed that the USB hub interfacing the particle counter and a few other devices to cominaux had no indicator lights on. I tried replacing its DC power supply, but still no response.
      • Finally, I bypassed the hub and plugged the particle counter directly into cominaux. After another 'systemctl restart rts-edc' on cymac1, the particle counts are now being recorded to frames. (note that only the particle counter temperature channel is recording; the AD590 temperature monitor is not being supplied with power)
Attachment 1: Screenshot_from_2022-04-20_14-41-11.png
Screenshot_from_2022-04-20_14-41-11.png
  2874   Tue Apr 19 14:50:00 2022 aaronNoise HuntingNoise Budgetphotodiode dark noise measurement

Things still look pretty weird. Today I'll measure the photodiode dark noise (the previous curve was theoretical, based on the known transimpedance of the 1811 and properties of the PDH electronics).

Measurement steps:

  1. Measure the PDH error signal with beam blocked, sending the output of the PDH box directly to SR785.
    • blocked beam with a razor blade dump
    • Disconnected the IF output from the PDH mixer/lowpass from the LB box input, and instead connected it to the SR785 input (1 MOhm)
    • Removed the RF power splitter and second PDH mixer from the system. These electronics are for locking both lasers to the cavity, but we want to understand the noise for a single laser locked to the cavity first. The electronics are still on the table and straightforward to set up.
    • I then recorded the power spectrum using labutils/netgpibdata/SRmeasureWideSP.py. The parameters .yml file is in cryo_lab/data/PD_dark (along with the measurement results).
      • The noise has a 1/f corner near 40 Hz, then is flat until about 20 kHz when it increases in a series of peaks. Since it's uncalibrated, I'll say no more on the overall noise level than that there is a discontinuity in the spectrum as recorded (possibly indicating a range issue for part of the measurement?).
      • When I went to turn off the lab electronics (below), I found that the OCXO preamplifier box was already off from when Shruti and I were rearranging some cables last time... so, today's measurement was a complete waste of time crying
  2. Calibrate the PDH error signal -- did not complete today
    1. Drive the laser current above 1 kHz and just enough frequency deviation that the full PDH error signal and both sideband resonances are visible on an oscilloscope trace
    2. Use [calibratePDH.ipynb?] to determine the slope of the PDH error signal near the DC resonance. The x-axis is converted from 'seconds' to 'Hz' by assuming the distance between the sideband resonances is twice the sideband frequency.

Misc

There will be a planned electrical outage tomorrow morning at 6 am. I turned off the following electronics before leaving the lab:

  • All laser diodes (2 Rio lasers on the PSOMA table)
  • All laser current and TEC drivers (2 Thorlabs drivers on the PSOMA rack, 2 custom drivers on the cryo cavs rack)
  • cominaux, spirous, gaston, and cymac1
  • moku pro
  • RF amplifier used for delay line measurement
  • both Sorensen DC supplies (after turning off the electronics they power)
  • Acromag chassis, and AA and AI chassis for the fast DAQ
  • Marconi, spectrum analyzers, SR560s, function generators, oscilloscopes, LB servo box
  • CCTV
  2873   Tue Apr 19 12:07:46 2022 shrutiLab InfrastructurePSOMAPSOMA table enclosure assembly

[Aaron, Shruti]

We managed to install one of the HEPA filter units on our enclosure today, the other one could not be installed because it did not have enough clearance of the overhead tubelight.

  2872   Fri Apr 15 16:25:00 2022 ranaNoise HuntingNoise Budgetlaser frequency noise measurement

The ITC 502 has an analog input modulation impedance of 10 kOhm, according to the manual. How did you set up the 16 dB attenuation?

  2871   Thu Apr 14 13:45:08 2022 aaronNoise HuntingNoise Budgetlaser frequency noise measurement

[shruti, aaron]

We are measuring the frequency noise on the beat note between the north and south lasers. The level of observed frequency noise on the beat note should tell us whether the calibration error is in our total measured frequency noise curve or current noise curve.

Setup

  • We used the Agilent 4395A to observe the North-South beat note and confirm we are sending 10 dBm to the delay line box
  • We then tuned the beat note to near 43 MHz and nulled the output of the delay line box
  • We are monitoring the (SR560, G=1 buffered) output of the delay line box using an oscilloscope and the Moku:Pro
    • With Moku:Pro's multi-instrument mode, we use moku's PID controller to apply feedback to the temperature tune port of the ITC510 driving the north laser and hold the beat note on the delay line null. (attachment 2)
    • Simultaneously, we record the power spectrum of the delay line box output. We saved the trace of the spectral density from the delay line output at several frequency ranges to produce the frequency noise curve for the beat note attached
    • To calibrate the delay line, we drive the ITC502 mod in port (20 mA/V) with 0.01 Vpp at 3.14 kHz passed through 16 dB attenuation. The delay line output has a 1.01 mVpp peak at the drive frequency. Assuming the laser diode response is 150 MHz/mA, this means the delay line responsivity is 4.7 MHz / mV.
      • Shruti also measured the delay line responsivity by introducing a small offset from null in the temperature control loop, and observing the difference in beat note frequency on the spectrum analyzer. She got a similar responsivity calibration (within a factor of 2) *

Unfortunately today's measurement did not agree with any of the previous measurements on the noise budget. We'll next try re-measuring the transfer functions and frequency noise of the locked cavity.

* Shruti: I added an offset to the digital PID control on the Moku:Pro at \pm0.5 mV from the locking point which was at -0.4 mV to zero the delay line using a simple Integrator with a UGF of 312 mHz. At the two points the peak of the beat note was at 44.43 MHz and 42.42 MHz respectively resulting in a calibration factor \frac{44.15-42.42}{1} = 1.73 \textrm{ MHz/mV} .

 

Attachment 1: CalibratedNoiseBudget.pdf
CalibratedNoiseBudget.pdf
Attachment 2: IMG_16928781279A-1.JPEG
IMG_16928781279A-1.JPEG
  2870   Wed Apr 13 11:10:28 2022 aaronNoise HuntingNoise BudgetPSOMA noise budget

New measurements

  • Current noise
    • Measuring current noise following the procedures in elog 2737.
      • I'm sensing the voltage across the laser diode from the DB9 breakout to an SR560 (G=100, AC-coupled, 100 kHz lowpass filter with 6 dB/oct rolloff) to an SR785
      • Updated the noise budget accordingly (see attachment 1), but must still have a calibration error because there's still a disagreement with the measured frequency noise
      • Not sure why the current noise turns up at high frequency for both ITC502 and ITC510
    • Diode sensing resistance
      • Near the operating point for the South laser diode (8 kOhm TEC resistance, 117 mA LD current), driving the ITC502 with 1 Vpp at its current modulation input results in .134 Vrms fluctuations across the laser diode as measured at the DB9 breakout by a fluke multimeter in AC voltage mode
        • The driver's modulation coefficient is 20 mA/V, which means the resistance of the South Rio laser diode near this operating point is 19 Ohms
      • Near the operating point for the North laser diode (7.5 kOhm TEC resistance, 143 mA LD current), dirving the ITC510 with 0.1 Vpp at its current modulation input results in 0.073 Vrms fluctuations
        • The driver's modulation coefficient is 100 mA/V, so the resistance of the N Rio laser diode near this operating point is 21 Ohms

misc

  • added new DB9 cable in fiber box, eliminating one DB9 adapter in the laser diode current drive path
Attachment 1: CalibratedNoiseBudget20220411.pdf
CalibratedNoiseBudget20220411.pdf
  2869   Tue Apr 12 15:00:39 2022 Ian MacMillanMiscEquipment LoanAccelerometer/seismometer loan Wilcoxon 731A

I borrowed a Wilcoxon Seismometer Model number 731A from the cryo lab. I am also going to borrow an amp/power source for it. Currently at the 40m

Attachment 1: IMG_1558.jpeg
IMG_1558.jpeg
  2868   Tue Apr 12 12:16:58 2022 AaronDailyProgressSi fabBasic cantilever scribe and break

This morning, I used the Dynatex scriber-breaker at the KNI to make three silicon cantilevers 7cm x 1cm x 300um. These are similar dimensions to those used in Zach's thesis. I used a fresh from the shipping box (no RCA, no PECVD) Boron doped 3" wafer. The wafer broke along the crystal axis while I was loading it into the scriber, and two of the 7cm cantilevers include a small section of the curved wafer edge. This should be acceptable for now, and could be improved by using 4" wafers.

The scribe and break process takes under an hour.

Attachment 1: 51649EDB-A7BC-49DC-AD53-F95345413D0E.jpeg
51649EDB-A7BC-49DC-AD53-F95345413D0E.jpeg
Attachment 2: 2741BA36-BF23-4289-9BBC-6054A10A3ECC.jpeg
2741BA36-BF23-4289-9BBC-6054A10A3ECC.jpeg
  2867   Mon Apr 11 16:30:12 2022 aaronNoise HuntingNoise BudgetPSOMA noise budget

We'd like to progressively improve the noise performance of the PSOMA cavity by updating a noise budget when we make changes. Attached is the noise budget we have now, which I'll be updating throughout the week. The budget contains rough estimates of the following contributions, and caveats or notes are listed below each source. Full calculations are on git.ligo.org in NoiseSpectraCalibration.ipynb

  • Frequency noise of the Rio lasers
    • Laser frequency noise is that estimated in elog 2807
    • This estimate only extends to about 100 Hz, since that was the target of the previous current noise measurement. We can of course repeat the three corner hat measurement, use the beat note between the two Rio lasers as a conservative estimate of the single-laser frequency noise, or find a different elog of Johannes' or Zach's that extends to higher frequency.
    • Given this, I've just used 20 kHz/rtHz at 1 Hz with a 1/f dependence extended in both directions for the Rio laser frequency noise
  • Photodiode dark noise
    • Photodiode dark noise is not measured, but calculated using known properties of the Newport 1811 photodiode: 2.5 pA/rtHz current noise across 40,000 Ohm transimpedance, resulting in 100 nV/rtHz voltage noise at the output of the photodiode.
    • The voltage noise out of the photodiode is passed through the PDH electronics, including a level 17 mixer and lowpass filter that together introduce up to 6 dB of loss. This results in 100 uV/rtHz noise on the PDH signal due to photodiode dark noise.
    • Finally, the voltage noise out of the PDH electronics is referred back to laser frequency noise using a cavity response of 123 mV/MHz and cavity pole at 2.26 MHz per elog 2846. The DC noise level due to photodiode dark noise is 800 Hz/rtHz, and rises like f^2 above the cavity pole.
  • Shot noise
    • I estimated that the incident laser power on the cavity is 200 uW and that we have 75% mode matching, but didn't locate a record of laser power at the time that the frequency noise was measured. I'm estimated the DC power incident on the photodiode with the cavity locked at 50 uW.
    • The responsivity of the photodiode at 1550 nm is about 1 A/W, and the noise spectral density from Schottky S=2e|I| (where e is the electron charge). The shot noise is propagated through the same PDH and cavity responses as the PD dark noise.
  • Current drive noise
    • Laser current noise is just wrong
    • The measured frequency noise is for the South laser diode driven by an ITC502. The current plotted is for the ITC510 driver as measured one year ago. The noise of the ITC510 is known to be higher than ITC502, but even then the reported current noise is too high to be believed. The current noise measurement is quick, so I'll repeat it this week and update that curve.
  • Seismic noise
    • Seismic noise is a very rough estimate. I used x=10^{-8}\mathrm{m/ } \sqrt{\mathrm{Hz}} * \frac{1+(f_\mathrm{cantilever}/f)^2}{f^2}
    • I'm not totally confident about this expression, but it's due to 1/f^2 dependence for ground motion highpassed through a cavity with a single suspended optic
  • Total of estimated noise contributions
    • I've ommitted the total noise curve at this time... I think it's not useful at this time as several of the noise conributions lie above the measured noise curve.
  • Measured frequency noise
    • The measured frequency noise is that reported in elog 2846

There are obviously major discrepancies between the measured noise and the estimated noise contributions. I'll work on correcting the errors in either the frequency noise measurement and calibration or the estimates this week.

Attachment 1: CalibratedNoiseBudget20220411.pdf
CalibratedNoiseBudget20220411.pdf
  2866   Wed Mar 30 16:18:58 2022 shrutiDailyProgressPSOMAMixing signal and pump (north and south)

I fixed the fiber connections, re-organized the fibers a little, and finally was able to see a little less than 1.5 mW of power contribution from each laser at the output of the fiber-to-free space coupler.

PM 50-50 BS convention also suggests that the diagram we had in the previous configuration was wrong.

North-> REV REFL (red)

South-> FWD IN (white)

50% output taken from FWD TRANS (white)

 

 

  2865   Tue Mar 29 16:03:40 2022 shrutiDailyProgressPSOMAMixing signal and pump (north and south)

[shruti, aaron]

Summary

Our next step is to try and lock both lasers to the PSOMA cantilever. For this we changed the setup to mix the two beams in fiber after modulating them at slightly different frequencies. We changed it from 'old setup' to 'new setup' shown in Attachment 1. When we turned on the south laser after making this change we only saw 4 uW when we expected 2 mW. Will troubleshoot this tomorrow.

 

Before changing the setup to the 'new setup' we also tested the configuration Aaron had it earlier in elog 2864 and saw that the power levels were what we expected: 

  • North laser power at 90% output: 4.8 mW
  • Power after the EOM in north path: 2.5 mW
  • 10% input->90% output of BS: 280 uW
  • Adding south EOM output to 90% input: 3.8 mW
  • Without north: 3.5 mW

The north EOM does not seem to be problematic.

 

Attachment 1: addpump_.pdf
addpump_.pdf
  2864   Thu Mar 10 13:43:20 2022 aaronMiscLasermode matching and slow servo, three corner hat revisited

I'd like to improve the mode matching. Since it takes a while, I'll also set up a slow servo on the TEC controller to hold the laser near resonance.

  • Slow temperature control loop
    • We already had a cable carrying the relevant fast signals to the AA chassis (PDH error signal, which I connected to the LB box error monitor, after the summing amplifier; the PDH control signal, which I connected to the LB box Aux output; and REFL MON and TRANS MON, both of which are monitored by an oscilloscope on the table before going to AA chassis).
    • The monitor channels are reaching the AA chassis, but not reflected in cymac. It's been a while and there's a TIM error, so I restarted the models by running 'rtreset' on cymac1. Afterwards, mostly green indicators on X1IOP_GDS_TP and X0DAQ_DETAIL... but white boxes on the GPS timing signal (attachment 1; has been this way for at least a year). I checked the simulink model and realized I was just driving the wrong port of the AA chassis; quick fix.
    • Eventually remembered how to run the temperature control loop:
      • My calc channels weren't calculating, not sure why. Needed to systemctl restart SoftIOC.service from cominaux (also restarted modbusIOC.service for good measure).
      • Set the PID parameters as follows... it would be nice to set these up to initialize automatically. All unlisted values (like the proportional gain constant) were left at 0 or are updating calc channels.
        • channel name value
          X1:OMA-SLD_TEMP_TIMESTEP 0.1
          X1:OMA-SLD_TEMP_TT 0.1
          X1:OMA-SLD_TEMP_TF 0.1
          X1:OMA-SLD_TEMP_KI 0.1
          X1:OMA-SLD_TEMP_EN 1
      • in a tmux session on spirou, go to scripts/temp_control and run the python script PIDLocker_cavcontrol.py with the config file PIDConfig_SLDTemp.ini
      • I also updated the epics definition of X1:OMA-ERC_MON_RATIO with the new 'dark' values of TRANS_MON and REFL_MON
  • mode matching
    • With the temperature control loop running, I roughly doubled the mode matching efficiency (as estimated by the ratio of transmitted / reflected light) by adjusting the input steering optics
  • Three corner hat: Since I couldn't diagnose the full N laser beam path, I decided to throw together a synchronous three corner hat measurement
    • connected a bunch of fibers. Cleaned all fiber tips before (re)connecting
      • Moved the 99-1% fiber BS from the cryo cavs table into the PSOMA fiber box
      • Ran a 10 m patch cable from cryo cavs table to the PSOMA rack and into the fiber box. Where the cable crosses the room, I put it in a cable wrap (same model as the one carrying fiber from the rack to the PSOMA table).
      • The patch cable takes the Teraxion beam directly from the laser to the 1% coupled port of the 99-1% BS
      • I disconnected the 50-50 BS mixing N and S PSOMA lasers from the 1611 beat note PD, and instead sent the N-S mixed output to the 99% coupled port of the 99-1 BS. The output of the 99-1% BS now carries PSOMA N, PSOMA S, and Teraxion lasers, and I sent this beam into the 1611 PD.
    • Set up the Moku Pro in multi-instrument mode with an oscilloscope, spectrum analyzer, and two phasemeters
    • Turned the lasers on one at a time and check that I'm not saturating the 1611 PD... I'm up to 8 mW (of 10 mW saturation), so I'm seeing some higher order RF beats in the spectrum, but close enough for today.
    • Used the Agilent spectrum analyzer to scan around the PSOMA N and S lasers until I get all three beat notes on the Moku's spectrum analyzer. I identify the beat notes by scanning N or S lasers and seeing which pair moves.
    • Set up one phasemeter to track the two of the beats, and the other phasemeter to track the third. I used a 10 kHz bandwidth to make sure the beat note doesn't move out of band during the measurement, but didn't play around with lower BW.
    • Set one phasemeter to acquire data at 19.1 kHz with a 15s delay, then quickly switch over to the second phasemeter and start its data acquisition at 19.1 kHz about 15 s later. It would be nice if one could start recording data on all instruments at once, and I'm not sure what the final data files will look like, hopefully I can sync them -- maybe the data logger instrument does this?
    • Exported the files. I'll check them out on my laptop at my convenience.
  • Returning system to previous state: Shruti requested I leave the cavity as before, so I'm returning it to the transfer function measuring state with only one laser in the system.
    • swapped a couple fibers
      • Turn off the Teraxion laser, but I'm leaving it connected to the fiber box for convenience. Also turn of PSOMA lasers temporarily, and return them to the previous temperature and current settings.
      • Replace the 99-1% combiner output with the 50-50 mixer output, so only the NS beat note is sent to the 1611
      • The beam sent to the optical table was the 90-10% combiner output containing the North and South beams after the EOM. I replaced this with the output of the S EOM (containing the S laser only, with sidebands)
    • Swapped a few electronics in the PDH loop
      • Removed the splitter and second PDH mixer box from the loop. The REFL PD's AC output once again goes directly to the S laser's PDH mixer box.
      • Replaced the Pomona filter box with a 10 dB attenuator at the LB box input
    • replaced the monitor channels
      • Moku Pro is monitoring pickoff of the amplified NS beat note from the 1611 on channel A, LB box error mon on channel B, current control signal on channel C, and the buffered delay line phasemeter output on channel D.
      • 1611 beat note is sent to an RF amplifier, then the delay line box
      • Turned on the 'clicking' oscilloscope
    • LB box returned to previous state: gain knob at 5.3, pure integrator mode
    • turned off the temperature control loop and set TEMP_TUNE to 0
    • Confirmed the cavity still locks and the NS beat note is still under 100 MHz
Attachment 1: Screenshot_from_2022-03-10_14-26-10.png
Screenshot_from_2022-03-10_14-26-10.png
  2863   Wed Mar 9 14:20:26 2022 aaronDailyProgressElectronicsupdates to fiber patch cables, realignment, N laser in path, relock
  • Replaced the patch cable carrying PSOMA South laser from the fiber box to the table with 2x P3-1550PM-FC-10 cables (one for S, one for N), and wrapped them in the protective cable cover. Clamped the cables with rubber on the table and fastened the flexible cable wrap to the optical table and enclosure. Then had to slightly realign the fiber launch until the 00 mode locks again.
    • Also replaced the 10 dB attenuator at the LB box input with a 6 dB attenuator, to make up for the 50-50 RF power splitter I introduced before the PDH mixers yesterday
  • Completed the N laser path: 90% of pickoff BS to N EOM, then to [10% path of a 90-10 fiber splitter... asking around for a 50-50 splitter, unless I'm missing something the 10% path doesn't even register
    • cleaned any fiber tips before connecting
    • clamped with rubber the EOM
    • Found a 90-10 splitter off the shelf, but we're out of spare fiber component mounting trays so I left it in the plastic packaging and secured with Kapton tape
  • Locking activities
    • Reducing the laser power (due to the RF power splitter, and the fiber splitter that mixes N and S beams) obviously pushes us to higher open loop gains... but we're close to the phase margin, so that doesn't come easily
    • I added Shruti's loop compensating filter box before the LB servo input, which allowed me to increase the LB servo gain knob to 6.2 (up from 5.3, on a log scale) before the loop oscillates... although, there's still not enough gain to completely suppress the 40 Hz oscillation from the cantilever's fundamental mode. I also removed the 6 dB attenuator and incidentally swapped from (-B) input to (+A) input on the LB box.
      • It's probably also worth improving the mode matching again. A couple times I locked on to a 1-0 mode.
      • Also played around with the LFGL but noticed no qualitative difference in the system (especially REFL DC) with LFGL off or on.
    • Plan is to use the Moku Lab as a PID controller lieu of a second LB servo box. However, turning on the N laser doesn't even register on the REFL photodiode... this is surprising, the PD should definitely be sensitive enough to see the N laser even at 10% of full power (should be 400 uW).
      • Is this an alignment issue? Wrong part of the laser's T-I curve? Something wrong earlier in one of the fiber components (seems likely)? I couldn't locate the Thorlabs power meter, so had trouble testing these. Has anyone borrowed it? Might just be looking in the wrong places, but usually it's on the optics table.
  • Misc
    • The second oscilloscope on the optics table was making an annoying switching sound, so I turned it off.
    •  
  2862   Tue Mar 8 13:45:20 2022 aaronDailyProgressElectronicssecond PDH box and EOM

Entered the lab around 1:30pm to measure the frequency noise of the PDH error signal with the cavity locked. But I found some things had changed. Seems the system was set up to measure transfer functions. And, someone started setting up a monitor in the enclosure. Cool, but what happened?

I made a second homodyne mixer box (using ZFM-3H-S+ followed by SLP-5). I split off (ZFRSC-123-S+) the REFL signal to send half the RF to this second box, and connected the oscillator channel 2 output to the new LO in. I also placed the our second EOM in the fiber box, but haven't connected the fibers. In this stage, I'd like to mix the two lasers in fiber to avoid aligning a second beam in free space... but would need a spare 50-50 beamsplitter to do so without rearranging the order of fiber components.

  2861   Thu Mar 3 13:09:25 2022 aaronNoise HuntingLasersome noise budgets maybe

I came to measure to measure the laser frequency noise budget and see what I can learn.

I was having trouble finding a lock despite the beamspot flashing. Eventually switched the PDH signal to the (-B) input of the LB box (was on +A), and the cavity immediately locked.

not much else, I figured out what was going on in the notebooks in scripts/calibrate_noise

  2860   Tue Mar 1 17:08:47 2022 shrutiLab InfrastructurePSOMAPSOMA table enclosure assembly

[Aaron, Shruti]

We installed all acrylic panels on the enclosure today [Attachment 1]. Doing this required adjusting nearly all the lower horizontal bars multiple times to get the doors to fit perfectly, and hammering down the rails.

The last part of the enclosure building is to install the HEPA filter units. We wanted to do that today but were not able to because the units are missing power cords (need to order).

Attachment 1: WorkNB_Y3Q2.pdf
WorkNB_Y3Q2.pdf
  2859   Wed Feb 23 17:23:35 2022 ChrisElectronicsPSOMAcurrent driver status

The current driver now in the EE shop for troubleshooting (chassis S1500267/PCB S1500270) looks to be a D1200719-v3 PCB that has been modded to v4. The mod to v4 is also documented in the DCC traveler.

I repeated Zach’s 2017 coarse current adjust knob response test with this unit driving a 25 Ω dummy load, and got slightly different results. There’s a bigger offset at the zero knob setting, and saturation sets in above a knob setting of ~7.

It may be that Zach followed through on his stated plan to mod the circuit with a voltage limiter, like the one added in v7 of the schematic.

Next steps are to check the board for signs of this possible mod, and see if the observed misbehavior can be reproduced.

Other mods in v6/v7 of the circuit address high frequency oscillations and crossover issues, and we may want to implement those as well.
 

Attachment 1: coarseknob.png
coarseknob.png
ELOG V3.1.3-